Show that the function defined as 𝑓(π‘₯) = { π‘₯, π‘₯ ∈ β„š

π‘₯ 2 , π‘₯ ∈ ℝ βˆ’ β„š

is continuous at 1 and discontinuous at 2.


1
Expert's answer
2021-09-16T07:41:17-0400

"f(x)=\begin{cases}
x ~~~~~ \text{if } x\in \mathbb{Q}\\
x^2 ~~~ \text{if } x \in \R-\mathbb{Q}
\end{cases}\\
\text{To show that } f\text{ is continuous at 1. We must show that } \\ \lim_{x\to 1} f(x)=f(1)\\
\text{Suppose } f(x)=x\\
\ \lim_{x\to 1} f(x)=1\\
\text{Suppose } f(x)=x^2\\
\ \lim_{x\to 1} f(x)=1
\text{ Since } f(x)=1, \text{ then } f \text{ is continuous at } 1.\\\\"

b.

"\text{Consider the neighborhood of 2}\\(2-\frac{1}{n}, 2+\frac{1}{n}) n\in \N\\
\text{Then the neighborhood contain an irrational number } x_n n\in \N. \text{ Since irrational number is dense in } \R.\\
i.e. x_n\in (2-\frac{1}{n}, 2+\frac{1}{n})\\
\implies 2-\frac{1}{n}<x_n<2+\frac{1}{n}\\
\implies -\frac{1}{n}<x_n - 2<\frac{1}{n}\\
\implies |x_n-2|<\frac{1}{n}\\
\implies x_n \to 2 \text{ as } n\to \infty\\
\text{Now, } f(x_n)=x_n^2 ~~\forall n\in \N \text{ as } x_n \text{ is irrational}\\
\therefore \lim_{n\to \infty}f(x_n)= \lim_{n\to \infty} x_n^2=(\lim_{n\to \infty} x_n)^2=2^2=4\neq f(2)=2\\
\text{Thus, } f \text{ is discontinuous at } 2"